2. IR aera l Cer) hat separate a Separa an alse bre hent vanishes atnPon
Answer #1

By Stone-Weierstrass theorem, {\bf a} must vanish at some point x_0\in X. Thus,

{\bf a}\subseteq \{f\in C(X):f(x_0)=0\}

Note that the subalgebra

A:= \{f\in C(X):f(x_0)=0\}

is closed in C(X). Since A\supseteq {{\bf a}} , we have A\supseteq {\overline{\bf a}} ; thus, in order to show that A= {\overline{\bf a}} , we just need to show A C .

Consider B={\bf a}+\mathbb R1 . If b_1=a_1+c_1 and (l are elements in B={\bf a}+\mathbb R1 , where a1, a2 E a and C1, C2 ER , then

b_1b_2=(a_1a_2+c_1a_2+c_2a_1)+c_1c_2

Since {\bf a} is an algebra, we have a_1a_2+c_1a_2+c_2a_1\in {\bf a} ; therefore,

b_1b_2=(a_1a_2+c_1a_2+c_2a_1)+c_1c_2\in {\bf a}+\mathbb R1

This shows that B={\bf a}+\mathbb R1 is closed under multiplication. Since the other defining conditions of a subalgebra in C(X) are immediate for B={\bf a}+\mathbb R1 , we conclude that it is a subalgebra in C(X). Since {\bf a}\subset B and {\bf a} separates points, so does B. Since \mathbb R1\subset B and R1 vanishes nowhere, we conclude that B={\bf a}+\mathbb R1 is nowhere vanishing subalgebra of C(X) that separates points, By Stone-Weierstrass we know that B = C(X) . But we know that {\overline{\bf a}}+\mathbb R1 is closed (because R1 is one-dimensional and {\overline{\bf a}} is closed), and since B\subseteq {\overline{\bf a}}+\mathbb R1 , we conclude

C(X)={\overline B}\subseteq {\overline{\bf a}}+\mathbb R1

showing that C(X)={\overline{\bf a}}+\mathbb R1 .

If possible, suppose that {\overline{\bf a}}\subsetneq \{f\in C(X):f(x_0)=0\} ; then there exists feC(x such that T0 1o but f\notin {\overline {\bf a}} . Since f\in C(X)={\overline{\bf a}}+\mathbb R1 , we know that there is some non-zero constant c\in\mathbb R such that f-c\in{\overline{\bf a}} . Since {\overline{\bf a}}\subsetneq \{f\in C(X):f(x_0)=0\} , this implies f(x_0)-c=0~\Rightarrow~f(x_0)=c . But T0 1o implies 0=f(x_0)=c, a contradiction. Hence,  {\overline{\bf a}}= \{f\in C(X):f(x_0)=0\}.